OpenStudy (anonymous):

Which of these solutions satisfies the Rolle’s Theorem for the function f(x)= x sq.rt. of (x+6) in [-6, 0]?

OpenStudy (anonymous):

\[f(x)=x\sqrt{x+6}\] for sure yes?

OpenStudy (anonymous):

yes (:

OpenStudy (anonymous):

\[f(-6)=f(0)=0\] so you can use rolle's theorem says there is an number c between -6 and for which \[f'(c)=\frac{3(c+4)}{2\sqrt{c+6}}=0\]

OpenStudy (anonymous):

your job is just to solve for c, which you can pretty much do in your head. looks like the solution is \[c=-4\]

OpenStudy (anonymous):

Omg! Thanks so much! It all makes sense now! (: